OpenStudy (nicole143):

Finding all zeros in a polynomial function: I need help with figure out how to do this. I thought I understood it but I can't seem to get down to the right answer. Problem: Find all zeros of the function y = x^3 - 3x^2 - 9x I'll put up what work I did, so that if anyone knows where I went wrong or how to fix it they can show me.. Thank you in advance!

OpenStudy (nicole143):

Sorry for the hand writing.

OpenStudy (campbell_st):

start by setting the polynomial to zero \[0 = x^3 -3x^2 -9x\] next take x out as a common factor \[0 = x(x^2 -3x - 9)\] so now you know one of the zeros since x is a factor x = 0 is a zero you'll need to quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] to find the other zeros of the quadratic where a = 1, b = -3 and c = -9 hope this helps.

OpenStudy (campbell_st):

you mistake is that the polynomial doesn't have a constant term. so you can't use the rational root theorem.

OpenStudy (nicole143):

Oh, thank you. I'll try and get it down to the right answer. Will you be on long enough to check it after or help me again if i get stuck?

OpenStudy (campbell_st):

yep thats fine

OpenStudy (nicole143):

I think I went wrong again :/ Am I to come up with a fraction?

OpenStudy (nicole143):

After that where do you go? Do you not use any coefficients when finding zeros? @campbell_st

OpenStudy (tkhunny):

Given a zero, F, there is a factor (x-F). You DID use the coefficients. What was it that you were putting in the Quadratic Formula? Let's work a little harder on getting the Quadratic Formula correct. 1) The 2a is the denominator for BOTH terms, not just the radical term. 2) The square root of a positive number does NOT produce "i". Give it another go.

OpenStudy (nicole143):

So it would just end in x = -3 +- the square root of 45/2 with 2 under the whole thing?

OpenStudy (tkhunny):

Yes, just as campbell_st has shown it, up above.

OpenStudy (nicole143):

Okay, what do you do after you have the quadratic? do you look at both the pictures I put up? I had thought that you did a bit of both but my teacher said not to use a calculated so I don't have that to graph with to find what to use in the synthetic division.

OpenStudy (campbell_st):

so the additional 2 zeros are \[x = \frac{3 + 3\sqrt{5}}{2}....and... x = \frac{3 - 3\sqrt{5}}{2}\] as well as x = 0

OpenStudy (nicole143):

How did you get to the 3 + 3 square root of 5?

OpenStudy (campbell_st):

well its simply a case of simplifying the radical \[\sqrt{45} = \sqrt{9 \times 5} = \sqrt{9} \times \sqrt{5} = 3\sqrt{5}\]

OpenStudy (nicole143):

Ohh I see. Okay, is that all you have to do?

OpenStudy (campbell_st):

thats it... you now have the 3 zeros... there are numerous ways of finding zeros... but the most important is 1. look and think.... straight away I saw a factor of x in every term... removing it gave me 1 zero and a quadratic... that I knew I could manage good luck.

OpenStudy (nicole143):

Okay, thank you very much. Could you please check to see that every step is there and that it is correct? @campbell_st

OpenStudy (campbell_st):

you value of b = -3 so the start of the quadratic formula is -b so you have - (-3) which becomes 3 that's the only mistake I saw

OpenStudy (nicole143):

Thank you.. Oh and I don't mean to be a bother but could you please check to see if one other problem I did is right? @campbell_st

OpenStudy (tkhunny):

You may also want to be a little more deliberate with your notation. \(-3^{2}\) is that \(-(3^{2})\;or\;(-3)^{2}\)? It's not always clear what is intended with that negative out in front. More parentheses solve the problem entirely.

OpenStudy (campbell_st):

you have an error in the division you multiplied the divisor by 2x^2 which means the 1st subtraction is incorrect... 2x^3 - 3x^2 - 2x^3 - 8x^2 -------------- 5x^2

OpenStudy (nicole143):

Don't you times the x by 2x^2 to get 2x^3?

OpenStudy (campbell_st):

thats correct... but its subtraction so 2x^3 - 2x^3 = 0 and -3x^2 - (-8x^2) = 5x^2

OpenStudy (campbell_st):

look at it this way \[2x^2(x -4) = 2x^3 - 8x^2\]

OpenStudy (campbell_st):

you forgot the minus is 8x^2...it was left as a positive by the look of things.

OpenStudy (nicole143):

Oh my goodness -_- I didn't see that it was times by a negative 4 to get -8. Thank you ha I mess up with those a bit:/

OpenStudy (campbell_st):

its just a case of practice and being careful.

OpenStudy (nicole143):

Is this all right now?

OpenStudy (nicole143):

@campbell_st

OpenStudy (campbell_st):

looks good... the quadratic quotient can be factored if necessary